# Promptlinecolor=black,linewidth=2pt, leftmargin=1cm,rightmargin=1cm

### v2

@@ 4.2 Quantitative Evaluation
We evaluate the performance of synthetic data generation models from two perspective: Privacy leakage by Distance to Closest Records (DCR) and Machine Learning Efficiency (MLE) (8; 33).
@@
MLE. To assess the utility of our synthetic data, we use it to train supervised model and test prediction accuracy on real data ($D_{\text{test}}$). The Adult data is used for a classification task, while the other three datasets are used for regression. For classification, we fit logistic regression model, random forest, and Support Vector Machine model, calculating F1 score. For regression, we fit linear regression and random forest models and calculate $R^{2}$. We also fit models using real data $D_{\text{train}}$ as a gold standard for the MLE for comparison. As a result, MALLM-GAN generated high-quality synthetic tabular data across multiple datasets and training data size, outperforming baselines (Table 2), specially with small training sizes ($N=100$). This indicates MALLM-GAN’s robustness to smaller sample sizes, unlike baselines that require more data. MALLM-GAN also outperformed baselines on both public and private datasets, suggesting it does not rely on the pre-trained LLM’s memorization.
Table 2: Benchmark results evaluated by MLE
@@
Adult (F1 score) Insurance ($R^{2}$) ATACH ($R^{2}$) ERICH ($R^{2}$)
N=100 Real data 0.79 0.73 0.09 -0.30
CTGAN $0.72\pm 0.01$ $-0.76\pm 0.37$ $-0.74\pm 0.20$ $-0.22\pm 0.19$
TVAE $0.75\pm 0.03$ $-0.46\pm 0.64$ $-0.19\pm 0.13$ $-0.67\pm 0.16$
Be-GReaT $0.76\pm 0.03$ $-0.10\pm 0.09$ $-0.92\pm 0.28$ $-0.84\pm 0.37$
MALLM-GAN (ours) $\mathbf{0.78\pm 0.03}$ $\mathbf{0.57\pm 0.08}$ $\mathbf{0.01\pm 0.14}$ $\mathbf{-0.13\pm 0.15}$
N=200 Real data 0.76 0.74 0.35 -0.08
CTGAN $0.67\pm 0.03$ $-0.02\pm 0.14$ $-0.12\pm 0.11$ $-0.19\pm 0.24$
TVAE $0.65\pm 0.00$ $0.13\pm 0.33$ $0.05\pm 0.11$ $-0.34\pm 0.07$
Be-GReaT $\mathbf{0.73\pm 0.01}$ $0.02\pm 0.13$ $-0.22\pm 0.08$ $-0.29\pm 0.10$
MALLM-GAN (ours) $\mathbf{0.73\pm 0.01}$ $\mathbf{0.74\pm 0.03}$ $\mathbf{0.14\pm 0.07}$ $\mathbf{-0.08\pm 0.06}$
N=400 Real data 0.74 0.82 0.47 0.21
CTGAN $0.58\pm 0.04$ $-0.12\pm 0.11$ $-0.11\pm 0.13$ $-0.11\pm 0.13$
TVAE $0.75\pm 0.03$ $0.09\pm 0.11$ $0.05\pm 0.10$ $-0.22\pm 0.06$
Be-GReaT $\mathbf{0.77\pm 0.03}$ $0.49\pm 0.07$ $0.14\pm 0.09$ $-0.09\pm 0.07$
MALLM-GAN (ours) $0.76\pm 0.06$ $\mathbf{0.66\pm 0.06}$ $\mathbf{0.17\pm 0.02}$ $\mathbf{0.04\pm 0.06}$
N=800 Real data 0.72 0.88 0.31 0.07
CTGAN $0.67\pm 0.00$ $-1.70\pm 0.29$ $-0.07\pm 0.04$ $-0.05\pm 0.02$
TVAE $\mathbf{0.73\pm 0.02}$ $0.70\pm 0.03$ $0.21\pm 0.06$ $-0.22\pm 0.11$
Be-GReaT $\mathbf{0.73\pm 0.01}$ $0.75\pm 0.04$ $\mathbf{0.33\pm 0.04}$ $-0.14\pm 0.06$
MALLM-GAN (ours) $0.72\pm 0.00$ $\mathbf{0.78\pm 0.02}$ $0.15\pm 0.02$ $\mathbf{0.05\pm 0.01}$
@@
DCR distributions. The DCR metric assesses the realism and diversity of synthetic data. It determines whether synthetic data points are too similar to the real data points which might indicate potential privacy concerns or too dissimilar which could affect the utility of the synthetic data. The DCR $d(\mathbf{x}_{\text{syn}},D_{\text{real}})=\min_{\mathbf{x}_{\text{real}}\in D_{\text{real}}}l_{1}\text{-norm}(\textbf{x}_{\text{syn}},\textbf{x}_{\text{real}}).$ Low DCR indicates that synthetic data are very close to real data points, implying a risk of data leakage and privacy concerns, as synthetic data too closely mimic the real data. We chose to use $l_{1}$-norm distance to measure the distance between two data points  (5). For the categorical variables, the distance is 1 if two categories are different, otherwise 0. As a result, MALLM-GAN achieved similar or higher DCR levels compared to baselines (Fig. 2), implying effective privacy protection without compromising MLE. Overall, MALLM-GAN demonstrated superior performance in generating synthetic data that maintains both privacy and utility.
Figure 2: DCR for the ATACH2 (N=100). MALLM-GAN achieved similar or higher level of DCR without compromising MLE.
@@ 4.3 Ablation study
Number $n$ of example in in-context few shot learning. Due to the LLM’s limited context length, we implemented a“batches in a batch” method to leverage all training data within these constraints (Section 3.2.1). We varied the number $n$ of examples and found $n=2$ to be optimal, achieving high DCR without compromising MLE (Supplement Section 2.4).
Causal structure and Optimization. To assess the impact of each component on overall performance, we examined the contribution of the causal structure in the data generation process $\theta$ and the LLM as an optimizer. We compared the full model, which includes both components, to a version without them, similar to CLLM (25) without post-processing data selection. The ablation study, conducted on one public dataset (Insurance) and one private dataset (ERICH), showed that incorporating the causal structure alone significantly improved the MLE compared to a model with only in-context few-shot learning (25). The LLM optimizer further enhanced $\theta$, achieving the highest MLE. Incorporating external knowledge into LLMs has been shown to significantly improve the quality of generated text, similar to retrieval-augmented generation (RAG) (16). Our approach shares this concept by incorporating a “knowledge” graph but optimizes the knowledge itself through adversarial optimization.
@@
Insurance ($R^{2}$) $-0.0172\pm 0.4593$ $0.5383\pm 0.0916$ $0.5721\pm 0.0791$
ERICH ($R^{2}$) $-0.3207\pm 0.1380$ $-0.1476\pm 0.1118$ $-0.1301\pm 0.1503$
@@ 4.4 Optimization trajectory of data generation process
A key advantage of MALLM-GAN is its transparent data generation process, described in natural text, allowing us to observe the evolution of the mechanism during adversarial optimization. We illustrate this with a trajectory example of the causal structure (Fig. 3) and task instruction (Table 4) using the Insurance dataset. Initially, the causal structure derived from heuristics had no edges (Fig. 3). Over iterative optimization, a stable causal structure emerged. The task instructions evolved to include specific details (Table 4). Our objective was to reduce the discriminator’s accuracy, and the score decreased over iterations, thanks to the optimized causal structure and task instruction.
Figure 3: Trajectory of causal structure in data generation process over adversarial optimization. Orange if the variables were used. Blue if not used in the causal structure.
@@ 4.6 Limitations
The proposed framework has several shortcomings. Firstly, due to the limited context length of the LLM, our model struggles with high-dimensional datasets with too many categorical variables, which make the context information lengthy and reduce the success rate of data generation. Additionally, mimicking the traditional GAN framework, it suffers from convergence issues, making it challenging to define the convergence of the prompt optimization strategy. While our model performs well with small sample sizes, showing better results than other baselines, the improvement diminishes with larger datasets. Moreover, the training and generation process is costly when dealing with large data volumes.
@@ 5 Conclusion
We propose a novel framework to generate synthetic tabular data by leveraging multi-agent LLMs to address the limited sample size issues which is prevalent in healthcare. Compared with other LLLM-based methods, we propose a in-context learning approach which does not require fine-tuning on LLM but still leverage the whole data. We leverage causal structure as data generation process and mimic a GAN architecture to optimize the data generation process. We demonstrate that our model can generate high quality synthetic data while preserving the privacy of real data. Moreover, compared with other black box models, our proposed work enables transparent data generation that allows domain experts to control the process.
@@ 2.3 Computing resource details
2.4 Number $n$ of example in in-context few shot learning
Given the limited context length that the LLM can understand, we proposed “batches in a batch” method to leverage all training data in limited context length in the generator LLM (Section 3.2.1). We varied the number $n$ of few-shot examples by $n=1,...5$ and measured the MLE (Fig. 5) and DCR distribution (Table 6) to find the optimal number $n$. As a result, the increasing number $n$ of examples did not always increase the MLE of synthetic data (Fig. 5) but decreased the DCR (Table 6), thus increasing privacy concerns. Instead, $n=2$ achieved sufficiently high DCR without compromising MLE. The MLE did not increase with more examples because the more examples will increase the context length and the generator LLM overlook some key context information. On the other hand, the DCR decreased with more examples because the generator LLM is more likely to stick to copy the provided examples. Interesting, the increasing number of examples does not affect the DCR of synthetic data generated from public dataset (Adult, Insurance).
[…]
@@ 2.4 Number $n$ of example in in-context few shot learning
Adult 5, 7, 8 6, 8, 11 6, 8, 11 6, 8, 10 6, 8, 10
Insurance 33, 87, 187 61, 145, 298 64, 144, 438 64, 159, 505 53, 126, 297
ATACH2 66, 81, 95 75, 88, 104 73, 83, 97 68, 81, 94 71, 79, 93
ERICH 66, 79, 92 69, 83, 100 65, 79, 94 67, 79, 94 62, 74, 85
Table 6: Number $n$ of examples and DCR. 25%, 50% (Median), 75% quantile.

### v3

@@ 4.2 Performance Evaluation
We evaluate the performance of synthetic data generation models from two perspectives: Privacy leakage by Distance to Closest Records (DCR) and Machine Learning Efficiency (MLE) (10; 42).
@@
MLE. To assess the utility of our synthetic data, we use it to train supervised model and test prediction accuracy on real data ($D_{\text{test}}$). The Adult data is used for a classification task, while the other three datasets are used for regression. For classification, we fit logistic regression model, random forest, and Support Vector Machine model, XGBoost Classifier, calculating F1 score. For regression, we fit linear regression, random forest, XGBoost Regressor and calculate $R^{2}$. For each model, we report the average of the best scores for each random seed. We also fit models using real data $D_{\text{train}}$ as a gold standard of the MLE for comparison. As a result, MALLM-GAN generated high-quality synthetic tabular data across multiple datasets and training data size, outperforming baselines (Table 2), specially with small training sizes ($N=100$). This indicates MALLM-GAN’s robustness to smaller sample sizes, unlike baselines that require more data. MALLM-GAN also outperformed baselines on both public and private datasets, suggesting it does not rely on the pre-trained LLM’s memorization.
Table 2: Benchmark MLE results over 5 datasets. Baseline results were obtained from training the supervised models directly on the real data. SMOTE* interpolates data within the training set, thus it gets higher accuracy by copying training data and compromising DCR.
@@
Adult ($F1$) Asia ($F1$) Insurance($R^{2}$) ATACH($R^{2}$) ERICH($R^{2}$)
N=100 Real data 0.86 0.83 0.82 0.26 -0.04
SMOTE* $0.78\pm 0.01$ $0.83\pm 0.00$ $0.80\pm 0.01$ $0.27\pm 0.03$ $-0.15\pm 0.13$
TabDDPM $0.75\pm 0.01$ $-5.26\pm 0.42$ $-0.99\pm 0.33$ $-0.19\pm 0.05$
CTGAN $0.66\pm 0.06$ $0.63\pm 0.19$ $-0.09\pm 0.11$ $-0.40\pm 0.21$ $-0.33\pm 0.11$
TVAE $0.67\pm 0.05$ $0.83\pm 0.01$ $0.39\pm 0.15$ $-0.01\pm 0.07$ $-0.11\pm 0.12$
Be-GReaT $0.71\pm 0.03$ $0.83\pm 0.00$ $0.54\pm 0.10$ $-0.25\pm 0.23$ $-0.38\pm 0.12$
MALLM-GAN $\mathbf{0.79\pm 0.02}$ $\mathbf{0.83\pm 0.00}$ $\mathbf{0.72\pm 0.00}$ $\mathbf{0.27\pm 0.07}$ $\mathbf{-0.03\pm 0.07}$
N=200 Real data 0.85 0.83 0.83 0.27 0.16
SMOTE* $0.78\pm 0.04$ $0.83\pm 0.00$ $0.79\pm 0.02$ $0.31\pm 0.04$ $0.05\pm 0.06$
TabDDPM $0.60\pm 0.15$ $0.56\pm 0.14$ $-0.55\pm 0.33$ $-0.30\pm 0.06$
CTGAN $0.61\pm 0.02$ $0.71\pm 0.10$ $-0.12\pm 0.08$ $-0.27\pm 0.05$ $-0.19\pm 0.10$
TVAE $0.67\pm 0.05$ $0.82\pm 0.01$ $0.62\pm 0.05$ $0.08\pm 0.06$ $-0.08\pm 0.07$
BeGReaT $0.69\pm 0.05$ $0.82\pm 0.00$ $\mathbf{0.72\pm 0.03}$ $0.16\pm 0.06$ $-0.18\pm 0.16$
MALLM-GAN $\mathbf{0.77\pm 0.03}$ $\mathbf{0.83\pm 0.01}$ $0.69\pm 0.04$ $\mathbf{0.28\pm 0.07}$ $\mathbf{0.02\pm 0.02}$
N=400 Real data 0.83 0.84 0.85 0.31 0.18
SMOTE* $0.85\pm 0.03$ $0.84\pm 0.00$ $0.83\pm 0.00$ $0.32\pm 0.02$ $0.07\pm 0.05$
TabDDPM $\mathbf{0.82\pm 0.03}$ $\mathbf{0.79\pm 0.03}$ $0.36\pm 0.02$ $\mathbf{0.09\pm 0.04}$
CTGAN $0.63\pm 0.02$ $0.59\pm 0.17$ $-0.18\pm 0.10$ $-0.08\pm 0.07$ $-0.24\pm 0.10$
TVAE $0.71\pm 0.07$ $0.71\pm 0.07$ $0.62\pm 0.05$ $0.16\pm 0.08$ $-0.19\pm 0.06$
Be-GReaT $0.79\pm 0.04$ $0.79\pm 0.00$ $0.72\pm 0.03$ $0.20\pm 0.06$ $-0.13\pm 0.07$
MALLM-GAN $0.79\pm 0.02$ $\mathbf{0.83\pm 0.00}$ $0.71\pm 0.03$ $\mathbf{0.27\pm 0.04}$ $0.02\pm 0.03$
N=800 Real data 0.71 0.84 0.85 0.40 0.21
SMOTE* $0.71\pm 0.03$ $0.84\pm 0.00$ $0.83\pm 0.00$ $0.37\pm 0.03$ $0.10\pm 0.05$
TabDDPM $0.70\pm 0.03$ $\mathbf{0.83\pm 0.01}$ $-0.53\pm 0.45$ $\mathbf{0.12\pm 0.04}$
CTGAN $0.64\pm 0.05$ $0.48\pm 0.06$ $-0.41\pm 0.06$ $-0.05\pm 0.06$ $-0.04\pm 0.02$
TVAE $0.77\pm 0.02$ $0.82\pm 0.01$ $0.68\pm 0.01$ $0.12\pm 0.07$ $-0.05\pm 0.03$
Be-GReaT $0.75\pm 0.07$ $0.82\pm 0.00$ $0.53\pm 0.21$ $0.00\pm 0.07$ $-0.04\pm 0.05$
MALLM-GAN $\mathbf{0.80\pm 0.02}$ $\mathbf{0.84\pm 0.00}$ $0.72\pm 0.01$ $\mathbf{0.36\pm 0.02}$ $0.02\pm 0.02$
@@
DCR distributions. The DCR metric assesses the realism and diversity of synthetic data. It determines whether synthetic data points are too similar to the real data points (potential privacy leakage) or too dissimilar (hurting the utility of the synthetic data). The DCR is defined as $d(\mathbf{x}_{\text{syn}},D_{\text{real}})=\min_{\mathbf{x}_{\text{real}}\in D_{\text{real}}}l_{1}\text{-norm}(\textbf{x}_{\text{syn}},\textbf{x}_{\text{real}}).$ Low DCR indicates that synthetic data are very close to real data points, implying a privacy leakage, as synthetic data too closely mimic the real data. We chose to use $l_{1}$-norm distance to measure the distance between two data points  (6). For the categorical variables, the distance is 1 if two categories are different; otherwise, 0. As a result, MALLM-GAN achieved similar or higher DCR levels compared to baseline models (Fig. 2), implying effective privacy protection without compromising MLE. Overall, MALLM-GAN demonstrated superior performance in generating synthetic data with small data as balancing privacy and utility.
Figure 2: DCR between the synthetic data and the real data. DCR were calculated based on training data and held-out test data for each model. A good model should have similar distributions between the DCR to training and the DCR to held-out dataset.
@@ 4.3 Ablation study
Number $n$ of example in in-context few shot learning. Due to the LLM’s limited context length, we implemented a“batches in a batch” method to leverage all training data within these constraints (Section 3.2.1). We varied the number $n$ of examples and found $n=1$ to be optimal, achieving high DCR without compromising MLE (Supplement Section 2.5).
Causal structure and Optimization. To assess the impact of each component on overall performance, we examined the contribution of the causal structure in the data generation process $\theta$ and the LLM as an optimizer. We compared the full model, which includes both components, to a version without them, similar to CLLM (31) without post-processing data selection (Table 3). The ablation study showed that incorporating the causal structure alone did not significantly improve the MLE compared to a model with only in-context few-shot learning. However, the LLM optimizer improved $\theta$ using prior knowledge encoded in LLM and finally achieved the highest MLE. Incorporating external knowledge into LLMs has been shown to significantly improve the quality of generated text, similar to retrieval-augmented generation (RAG) (20). Our approach shares this concept by incorporating a “knowledge” graph but optimizes the knowledge itself through adversarial optimization.
@@
Adult ($F1$) $0.7550\pm 0.0454$ $0.7503\pm 0.0393$ $\mathbf{0.7892\pm 0.0358}$
Asia ($F1$) $0.2335\pm 0.0000$ $0.2756\pm 0.2842$ $\mathbf{0.8282\pm 0.0041}$
Insurance ($R^{2}$) $0.6821\pm 0.0193$ $0.6718\pm 0.0916$ $\mathbf{0.7152\pm 0.0447}$
ATACH ($R^{2}$) $0.1581\pm 0.0850$ $0.1326\pm 0.0637$ $\mathbf{0.2726\pm 0.0707}$
ERICH ($R^{2}$) $-0.0647\pm 0.0701$ $\mathbf{0.0281\pm 0.0424}$ $-0.0253\pm 0.0671$
@@ 4.4 Optimization trajectory of data generation process
A key advantage of MALLM-GAN is its transparent data generation process, described in natural text, which allows us to observe the evolution trajectory of the data generation mechanism during adversarial optimization. We present examples of optimization trajectories. We showed how the causal structure evolves to ground truth (Fig. 3) over iteration. We used the Asia dataset because it has known ground-truth causal structures and reported graph edit distance (GED) between ground truth and identified causal structures. In this example, the heuristically initialized causal structure gradually converges to ground truth, thanks to the knowledge obtained from the pre-trained LLM. Different convergence patterns were observed with different initialization strategies (Table 9), supporting the benefit of our heuristic initialization. We also investigated how the task instruction in the generator prompt gets sophisticated and how the discriminator’s accuracy changes over iterations. In Table 4, the task instructions evolved to include specific details, and the discriminator’s accuracy decreased, implying that synthetic data gets indiscriminative to real data.
Figure 3: An example of trajectory of causal structure in data generation process over adversarial optimization using Asia dataset. T: Tuberculosis, V: Visit to Asia, S: Smoke, LC: Lung cancer, T/L: Tuberculosis or Lung cancer, CX: Chest X-ray, D: Dyspnea, B: Bronchitis
@@
Epoch 1 “The ultimate goal is to produce accurate and convincing synthetic data that dutifully represents these causal relationships given the user provided samples.” 100.0%
Epoch 2 “The ultimate goal is to create a detailed and convincing dataset that accurately mirrors these causal pathways. While synthesizing your data, keep in mind the following key relationships: a ’visit to Asia’ increases the likelihood of ’tuberculosis’, ’smoking’ can lead to ’lung cancer’ and ’bronchitis’, and both ’tuberculosis’ and ’lung cancer’ can contribute to ’either tuberculosis or lung cancer’, which in turn can lead to ’Dyspnea’. Also, take note of how both ’tuberculosis’ and ’lung cancer’ are associated with ’chest X-ray’ results. Your data should reflect these intricate relationships while remaining consistent and realistic.” 76.19%
Epoch 4 “You are tasked with generating a synthetic dataset that faithfully demonstrates the given causal connections. Make sure the dataset illustrates how a ’visit to Asia’ can cause ’tuberculosis’, how ’smoking’ can lead to ’lung cancer’ and ’bronchitis’, and how either ’tuberculosis’ or ’lung cancer’ can eventually incite ’Dyspnea’. Also, the dataset should reasonably reveal how a ’chest X-ray’ ties in with ’tuberculosis’ and ’lung cancer’. Ensure the synthetic data reflects realistic scenarios where these factors interact, affecting each other exactly as per these defined causal relationships.” 66.67%
@@ 4.6 Limitations
The proposed framework has several shortcomings. Firstly, due to the limited context length of the LLM, our model struggles with high-dimensional datasets with too many categorical variables, which make the context information lengthy and reduce the success rate of data generation. Another limitation of synthetic data generation introduced by LLM is that the LLM struggles with random number generation as pointed out in (15), which cast negative effects on our framework’s potential when dealing with datasets of many continuous variables. Additionally, mimicking the traditional GAN framework, it suffers from a theoretical convergence guarantee. While our model performs well with small sample sizes, showing better results than other baselines, the improvement diminishes with larger datasets. Moreover, the training and generation process is costly when dealing with large data volumes.
@@ 5 Conclusion
We propose a novel framework to generate synthetic tabular data by leveraging multi-agent LLMs to address the limited sample size issues that are prevalent in healthcare. Compared with other LLM-based methods, we propose an in-context learning approach that does not require fine-tuning on LLM but still leverages the whole data. We use causal structure to guide the data generation process and mimic a GAN architecture to optimize the process. We demonstrate that our model can generate high-quality synthetic data while preserving the privacy of real data. Moreover, compared with other black box models, our proposed work enables transparent data generation that allows domain experts to control the process.
@@ 2.1 Benchmark datasets descriptions
Asia 10000 8 This is the dataset used to illustrate the utility of Baysian network to do causal structure discovery. The dataset is available in the R-package(30).
@@ 2.2 Hyperparameters
  • TabDDPM: Default parameters
2.3 Comparison among different kinds of discriminators
Table 6: Comparison of different discriminators effects on the quality of the synthetic data. An experiments on sub-sample of Adult data.
N = 100 N = 200 N = 400 N = 800
Adult (F1 score) XGBoost $0.78\pm 0.03$ $0.73\pm 0.01$ $0.76\pm 0.06$ $0.72\pm 0.00$
Logistic regression $0.79\pm 0.02$ $\mathbf{0.77\pm 0.02}$ $\mathbf{0.79\pm 0.03}$ $\mathbf{0.80\pm 0.02}$
Neural Network $\mathbf{0.80\pm 0.02}$ $0.57\pm 0.12$ $0.78\pm 0.06$ $0.67\pm 0.12$
@@ 2.4 Computing resource details
2.5 Number of examples in in-context few shot learning
Given the limited context length that the LLM can understand, we proposed “batches in a batch” method to leverage all training data in limited context length in the generator LLM (Section 3.2.1). We varied the number $n$ of few-shot examples by $n=1,...5$ and measured the MLE (Fig. 5) and DCR distribution (Table 7, 8) to find the optimal number $n$. As a result, the increasing number $n$ of examples did not always increase the MLE of synthetic data (Fig. 5) but decreased the DCR (Table 7), thus increasing privacy concerns. Instead, $n=1$ achieved sufficiently high DCR without compromising MLE. The MLE did not increase with more examples because the more examples will increase the context length and the generator LLM overlook some key context information. On the other hand, the DCR decreased with more examples because the generator LLM is more likely to stick to copy the provided examples. Interesting, the increasing number of examples does not affect the DCR of synthetic data generated from public dataset (Adult, Insurance).
Figure 5: Number $n$ of examples and MLE.
@@ 2.5 Number of examples in in-context few shot learning
Adult 5, 6, 10 5, 7, 12 4, 6, 9 4, 6, 10 4, 6, 11
Insurance 31, 93, 301 44, 66, 453 32,60, 182 33, 73, 167 29, 55, 168
Asia 0, 0, 0 0, 0, 0 0, 0, 0 0, 0, 0 0, 0, 0
ATACH2 61, 73, 88 72, 87, 97 69, 78, 89 66, 75, 94 67, 83, 104
ERICH 53, 61, 79 59, 78, 96 59, 74, 101 56, 72, 96 50, 64, 88
Table 7: Number $n$ of examples and DCR to training dataset. 25%, 50% (Median), 75% quantile.
Table 8: Number $n$ of examples and DCR to held out dataset. 25%, 50% (Median), 75% quantile.
1 2 3 4 5
Adult 4, 7, 10 5, 7, 11 5, 6, 10 4, 7, 10 4, 7, 11
Insurance 30, 115, 337 34, 91, 405 36, 76, 245 24, 64, 170 27, 70, 150
Asia 0, 0, 0 0, 0, 0 0, 0, 0 0, 0, 0 0, 0, 0
ATACH2 84, 100, 120 82, 99, 122 81, 97, 125 79, 98, 124 82, 103, 128
ERICH 70, 87, 110 66, 82, 111 51, 82, 104 62, 80, 108 62, 80, 117
2.6 Optimization trajectory: example on insurance dataset
As seen in Table 9, the causal structure does not converge to the ground truth after 5 epochs. When initialized with the ground truth, the causal structure maintains slight fluctuations. Another example of causal structure discovery on insurance dataset were presented in Supplementary 2.6.
Table 9: Graph Edit Distance (GED) between causal structure in $\theta$ and ground truth
Epoch No Initialization Heuristic Initialization Ground Truth Initialization
0 16 5 0
1 4 1 0
2 4 1 0
3 3 0 0
4 4 0 2
Here is the example for the Insurance dataset. Initially, the causal structure derived from heuristics had no edges (Fig. 6). Over iterative optimization, a stable causal structure emerged. The task instructions evolved to include specific details (Table 10). Our objective was to reduce the discriminator’s accuracy, and the score decreased over iterations, thanks to the optimized causal structure and task instruction.
Figure 6: Trajectory of causal structure in data generation process over adversarial optimization.
Iteration Task instruction Score